OpenStudy (anonymous):

I am having to do a assignment over that i did too long ago to really remember. That and the fact that I never was good at these anyways. But here i am expected to show my work (: To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem.

OpenStudy (amistre64):

m + f = total m+ 5f = 25 m+10f = 40

OpenStudy (anonymous):

subtract 5f=15 f=3

OpenStudy (amistre64):

subtract the m's like this: m+ 5f = 25 m+10f = 40 (-1) m+ 5f = 25 -m -10f = -40 ------------- -5f = -15 f = 3 m+5(3) = 25 m+15 = 25 m = 10 ; dbl check 10 + 10(3) = 40 10 + 30 = 40

OpenStudy (anonymous):

Hmmm alright amistre (: Thanks a BUNCH! And seriously? 407 fans and HOW many medals? Jeezus!

OpenStudy (amistre64):

lol. I have a blacksmith shop in the back; and cant figure out the fanage tho lol

OpenStudy (anonymous):

Hahahaha Always the joker (:

OpenStudy (amistre64):

but i get to see the kids this weekend; and go to normal college in the fall; so ita all good ;)

OpenStudy (anonymous):

thanks a lot ivy

OpenStudy (anonymous):

Welcome (: